OpenStudy (anonymous):

A ball of mass 0.20 kg is thrown vertically up with an initial velocity of 20m/s . Calculate the max. PE it gains as it goes up.

OpenStudy (anonymous):

@sauravshakya help ?

OpenStudy (anonymous):

FIND THE INITIAL KINETIC ENERGY

OpenStudy (anonymous):

CAN U?

OpenStudy (anonymous):

yeah 40 is right ?

OpenStudy (anonymous):

K.E=1/2 mv^2 =1/2*0.20*(20)^2=40 J

OpenStudy (anonymous):

is that all i have i to do ?

OpenStudy (anonymous):

So, that will be the max.PE it gains

OpenStudy (anonymous):

okay , ty :)

OpenStudy (anonymous):

WELCOME